# MultiModalPFN: Extending Prior-Data Fitted Networks

### v2

@@ 5 Conclusion
We introduced MMPFN, a multimodal extension of TabPFN that unifies tabular, image, and text inputs with per-modality encoders, a modality projector, and the TabPFN backbone. We developed MGM and CAS, which map non-tabular embeddings to the tabular space and mitigate token-count–induced attention imbalance. By leveraging pretrained foundation models and fine-tuning lightweight components, MMPFN achieved strong accuracy with substantially lower training costs. Across medical and general-purpose benchmarks, it consistently outperformed competitive state-of-the-art methods, scaled positively as modalities were added, and maintained robust performance in low-data regimes.
@@
Fig. S1 illustrates the relationship between the number of non-tabular input features generated by MGM and CAP and the resulting performance across the evaluated datasets. In all cases, the best performance is achieved when the number of non-tabular features is similar to the number of tabular features. Performance tends to degrade when the number of non-tabular features is either substantially smaller or larger than the number of tabular features. These results experimentally support our analysis of attention imbalance.
@@
Replacing Modality Encoder
@@
Table S2: Effect of replacing the image encoder. Performance of MMPFN when substituting DINOv2 with DINOv3. Results are reported as averaged accuracy over five random seeds.
@@
Figure S2: Effect of distribution-alignment and orthogonality regularization. Under the same experimental conditions as Fig. 2, we compare the performance of variants that incorporate orthogonality and MMD-based regularization constraints to the MGM baseline, for both image-only and tabular+image settings.
@@
Comparison with Patch-Token Features
@@
Empirically, we also observe that models using ViT patch-token outputs underperform those relying on the $[CLS]$-driven MGM features. On PAD-UFES-20, replacing the $[CLS]$-based MGM heads with patch-token features decreases accuracy by $0.85$ percentage points (from $85.22\%$ to $84.02\%$). We attribute this degradation to the fact that the $[CLS]$ representation of a well-trained foundation model already encodes task-relevant global information, while raw token-level outputs contain substantial redundancy and noise. Consequently, patch-token features are less suitable than $[CLS]$-based MGM heads for constructing compact, tabular-like feature representations.
Appendix S2 Experiments Setup
@@
For the training procedure, we adopt the official TabPFN repository [^2] https://github.com/LennartPurucker/finetune_tabpfn_v2 and modify it to support MMPFN by extending the MMPFNClassifier and MMPFNRegressor classes. Fine-tuning is performed by splitting the available data into training and validation sets and updating model parameters based on the validation loss. We use a small learning rate of $1\times 10^{-5}$, a fixed budget of $100$ training steps, and ScheduleFree [13] for learning rate scheduling. Fig. S3 shows the learning curve on four different datasets. For contrastive-pretraining baselines [23, 17, 16], we train for $500$ epochs using a cosine-annealing scheduler with a $10$-epoch warmup.
@@
We provide the detailed procedure used for the cosine-similarity–based correlation analysis in Figs. 3 and 4.3. The TabPFN encoder for tabular data normally groups multiple features into a single embedding to reduce memory usage, but such grouping can introduce noise when comparing cosine similarity between tabular and image (or text) embeddings. To obtain more precise relationships, we set the tabular group size to $1$, generate an individual embedding for each feature, and then compare these feature-wise embeddings with image embeddings.

### v3

@@ 5 Conclusion
We introduced MMPFN, a multimodal extension of TabPFN that unifies tabular, image, and text inputs with per-modality encoders, a modality projector, and the TabPFN backbone. We developed MGM and CAP, which map non-tabular embeddings to the tabular space and mitigate token-count–induced attention imbalance. By leveraging pretrained foundation models and fine-tuning lightweight components, MMPFN achieved strong accuracy with substantially lower training costs. Across medical and general-purpose benchmarks, it consistently outperformed competitive state-of-the-art methods, scaled positively as modalities were added, and maintained robust performance in low-data regimes.
Acknowledgements
This work was supported in part by the National Research Foundation of Korea (NRF) funded by the Korean government under Grants RS-2023-00221365 and RS-2024-00352566.
@@ Acknowledgements
Figure S1 illustrates the relationship between the number of non-tabular input features generated by MGM and CAP and the resulting performance across the evaluated datasets. In all cases, the best performance is achieved when the number of non-tabular features is similar to the number of tabular features. Performance tends to degrade when the number of non-tabular features is either substantially smaller or larger than the number of tabular features. These results experimentally support our analysis of attention imbalance.
@@
Parameter budgets in Table 5.
In Table 5, the parameter counts of the modality projector vary depending on the architectural choice. Let NN denote the number of MGM heads and $d$ the token dimension. The parameter counts scale as follows: single-head + linear $O(d)$, single-head + MLP $O(d^{2}+d)$, multi-head + MLP $O\!\left(N(d^{2}+d)\right)$, and multi-head + MGM $O\!\left(N(d^{2}+d/2)\right)$. MGM requires fewer parameters than the multi-head MLP due to its linear gating with channel splitting. These results indicate that the performance trends in Table 5 are not solely explained by increased model capacity. Moreover, the additional parameters introduced by the projector contribute only marginally to inference latency.
@@
Replacing Modality Encoders.
@@
We also examine the effect of replacing the text encoder. As shown in Table S3, switching between ELECTRA and DeBERTa results in only minor performance differences across the evaluated datasets. These results suggest that while stronger encoders can provide modest improvements, the overall performance of MMPFN remains relatively stable across different encoder choices.
Table S2: Effect of replacing the image encoder. Performance of MMPFN when substituting DINOv2 with ResNet50 and DINOv3. Results are reported as averaged accuracy over five random seeds.
@@
Table S3: Effect of replacing the text encoder. Performance of MMPFN when using different pretrained text encoders. Results are reported as averaged accuracy over five random seeds.
Encoder Airbnb Salary
Electra 47.78 46.17
DeBERTa 47.82 45.69
Figure S2: Additional results of token-count and attention mass. We extend the analysis in Figure 3 to additional datasets. For Mass and Calc, each sample contains three images, and thus the number of non-tabular tokens is given by the number of mgm heads times 3. The number of tabular tokens is 25, 4 and 4 for Airbnb, Mass, and Calc, respectively.
@@
Figure S3: Effect of distribution-alignment and orthogonality regularization. Under the same experimental conditions as Figure 2, we compare the performance of variants that incorporate orthogonality and MMD-based regularization constraints to the MGM baseline, for both image-only and tabular+image settings.
@@
Comparison with Patch-Token Features.
@@
Empirically, we also observe that models using ViT patch-token outputs underperform those relying on the $[CLS]$-driven MGM features. On PU20, replacing the $[CLS]$-based MGM heads with patch-token features decreases accuracy by $0.85$ percentage points (from $85.22\%$ to $84.02\%$). We attribute this degradation to the fact that the $[CLS]$ representation of a well-trained foundation model already encodes task-relevant global information, while raw token-level outputs contain substantial redundancy and noise. Consequently, patch-token features are less suitable than $[CLS]$-based MGM heads for constructing compact, tabular-like feature representations.
Additional qualitative results of attention imbalance.
To complement the analysis in the main paper, we further examine the relationship between token count and attention mass on additional datasets. As shown in Figure S2, we observe a consistent trend across all datasets: as the number of non-tabular tokens increases, the attention mass assigned to the non-tabular modality increases monotonically, while the attention to tabular tokens decreases accordingly.
For datasets such as Mass and Calc, each sample contains multiple images, and the number of non-tabular tokens is given by the number of MGM heads multiplied by the number of images. Despite variations in the number of tabular tokens across datasets, such as 25 for Airbnb and 4 for Mass and Calc, the same qualitative behavior consistently emerges. These results further indicate that attention allocation is primarily determined by the relative proportion of tokens within the input sequence.
@@
For the training procedure, we adopt the official TabPFN repository [^2] https://github.com/LennartPurucker/finetune_tabpfn_v2 and modify it to support MMPFN by extending the MMPFNClassifier and MMPFNRegressor classes. Fine-tuning is performed by splitting the available data into training and validation sets and updating model parameters based on the validation loss. We use a small learning rate of $1\times 10^{-5}$, a fixed budget of $100$ training steps, and ScheduleFree [11] for learning rate scheduling. Figure S4 shows the learning curve on four different datasets. For contrastive-pretraining baselines [20, 14, 13], we train for $500$ epochs using a cosine-annealing scheduler with a $10$-epoch warmup.
@@
We provide the detailed procedure used for the cosine-similarity–based correlation analysis in Figures 4 and 4.3. The TabPFN encoder for tabular data normally groups multiple features into a single embedding to reduce memory usage, but such grouping can introduce noise when comparing cosine similarity between tabular and image (or text) embeddings. To obtain more precise relationships, we set the tabular group size to 11, generate an individual embedding for each feature, and then compare these feature-wise embeddings with image embeddings.